Mary is five less than three times as old as sue. If the sum of their ages is 23, how old is mary?
Answer by stanbon(75887)

Mary is five less than three times as old as sue. If the sum of their ages is 23, how old is mary?
------
Equations:
m + s = 24
m = 3s-5
---
Substitute for "m" and solvd for "s":
3s-5 + s = 24
4s = 29
Sue = (29/4) yrs
---------------------
Solve for "m":
m + (29/4) = 24
----
4m + 29 = 96
4m = 67
Mary = (67/4) years
====================
Cheers,
Stan H.
============